OpenStudy (anonymous):

To calculate the enthalpy change for the reaction, 2CO (g) + O2 (g) → 2 CO2 (g), you can use ΔHf° values for each reactant and product. Which setup shows how the enthalpy change should be calculated? ΔH°reaction = [ΔHf° CO2(g)] – [ΔHf° CO(g) + ΔHf° O2(g)] ΔH°reaction = [2ΔHf° CO2(g)] – [2Δ Hf° CO(g) + ΔHf° O2(g)] ΔH°reaction = [2ΔHf° CO(g) + ΔHf° O2(g)] – [2ΔHf° CO2(g)] ΔH°reaction = [2ΔHf° CO(g) – ΔHf° O2(g)] + [2ΔHf° CO2(g)]

OpenStudy (freckles):

well according to my research just now reactants are before the arrow and products are the terms that follow the arrow and I looked up the formula and the formula is: \[\Delta H^o=\sum n H^o_{f_{ \text{products}}}- \sum m H^o_{f_{ \text{reactants}}}\] if that helps

OpenStudy (freckles):

still trying to figure out the n and m and the H type symbol there in the formula